3

message

If a sample of gas at a temperature of 24,2 °C, has a volume of 536 mL at 637 torr, what will its volume be if the pressure is increased to 712 torr?


1
Expert's answer
2021-08-29T16:56:52-0400

"\frac{pV}T=const,"

"pV=const,"

"V\sim \frac 1p,"

"V_2=V_1\frac{p_1}{p_2},"

"V_2=536\cdot \frac{637}{712}=480~ml."